# Debiased Machine Learning and Network Cohesion for Doubly-Robust Differential Reward Models in Contextual Bandits

### v2

@@ Abstract
A common approach to learning mobile health (mHealth) intervention policies is linear Thompson sampling. Two desirable mHealth policy features are (1) pooling information across individuals and time and (2) incorporating a time-varying baseline reward. Previous approaches pooled information across individuals but not time, failing to capture trends in treatment effects over time. In addition, these approaches did not explicitly model the baseline reward, which limited the ability to precisely estimate the parameters in the differential reward model. In this paper, we propose a novel Thompson sampling algorithm, termed “DML-TS-NNR” that leverages (1) nearest-neighbors to efficiently pool information on the differential reward function across users and time and (2) the Double Machine Learning (DML) framework to explicitly model baseline rewards and stay agnostic to the supervised learning algorithms used. By explicitly modeling baseline rewards, we obtain smaller confidence sets for the differential reward parameters. We offer theoretical guarantees on the pseudo-regret, which are supported by empirical results. Importantly, the DML-TS-NNR algorithm demonstrates robustness to potential misspecifications in the baseline reward model.
@@ 1 Introduction
Mobile health (mHealth) and contextual bandit algorithms share a connection in the realm of personalized healthcare interventions. mHealth leverages mobile devices to deliver health-related services for real-time monitoring and intervention. Contextual bandit algorithms, on the other hand, are a class of machine learning techniques designed to optimize decision-making in situations where actions have contextual dependencies. The synergy arises when mHealth applications deploy contextual bandit algorithms to tailor interventions based on individual health data and context. For example, in a mobile health setting, a contextual bandit algorithm might dynamically adapt the type and timing of health-related notifications or interventions based on the user’s current health status, historical behavior, and contextual factors like location or time of day Tewari & Murphy (2017).
At each decision point, a learner receives a context, chooses an action, and observes a reward. The goal is to maximize the expected cumulative reward. High-quality bandit algorithms achieve rewards comparable to those of an optimal policy. To achieve near-optimal performance in mHealth, bandit algorithms must account for (1) the time-varying nature of the outcome variable, (2) nonlinear relationships between states and outcomes, (3) the potential for intervention efficacy to change over time (due, for instance, to habituation as in  Psihogios et al. (2019)), and (4) the fact that similar participants tend to respond similarly to interventions (Künzler et al., 2019).
Traditional mHealth intervention development—including just-in-time adaptive interventions (JITAIs), which aim to tailor the timing and content of notifications to maximize treatment effect (Nahum-Shani et al., 2018)—has centered on treatment policies pre-defined at baseline (e.g., Battalio et al. (2021); Nahum-Shani et al. (2021); Bidargaddi et al. (2018); Klasnja et al. (2019)). As the development of JITAIs shifts towards online learning (e.g., Trella et al. (2022); Liao et al. (2020); Aguilera et al. (2020)), we have the opportunity to incorporate the four key characteristics listed above into the development of optimal treatment policies through algorithms such as contextual bandits.
Although some solutions to these problems have been presented, no existing method offers a comprehensive solution that simultaneously addresses all four challenges in a satisfactory manner. The purpose of this paper is to fill this gap with a method that performs well in the mHealth setting where data is high-dimensional, highly structured, and often exhibits complex nonlinear relationships. To that end, this paper offers three main contributions: (1) A novel algorithm, termed as “DML-TS-NNR” that flexibly models the baseline reward via the double machine learning (DML) framework and pools efficiently across both users and time via nearest-neighbor regularization; (2) theoretical results showing that DML-TS-NNR achieves reduced confidence set sizes and an improved regret bound relative to existing methods; and (3) empirical analysis demonstrating the superior performance of DML-TS-NNR relative to existing methods in simulation and two recent mHealth studies.
The paper proceeds as follows. Section 2 summarizes related work. Section 3 describes the model and problem statement. Section 4 describes the algorithm along with the resulting theoretical results. Section 5 describes experimental results for simulations and two mHealth studies. Section 6 concludes with a discussion of limitations and future work.
@@ 2 Related Work
The closest works are Choi et al. (2022) and Tomkins et al. (2021). Choi et al. (2022) employs a semi-parametric reward model for individual users and a penalty term based on the random-walk normalized graph Laplacian. However, limited information is provided regarding the explicit estimation of baseline rewards and the pooling of information across time. In contrast, Tomkins et al. (2021) carefully handles the issue of pooling information across users and time in longitudinal settings, but their approach (intelligentpooling) requires the baseline rewards to be linear and does not leverage network information. Below, we provide a summary of other relevant work in this area.
Thompson Sampling. Abeille & Lazaric (2017) showed that Thompson Sampling (TS) can be posed as a generic randomized algorithm constructed on the regularized least-squares (RLS) estimate rather than one sampling from a Bayesian posterior. At each step $t$, TS samples a perturbed parameter, where the additive perturbation is distributed so that TS explores enough (anti-concentration) but not too much (concentration). Any distribution satisfying these two conditions introduces the right amount of randomness to achieve the desired regret without actually satisfying any Bayesian assumption. We use the high-level proof strategy of Abeille & Lazaric (2017) in this work to derive our regret bound, although we need additional tools to handle our longitudinal setting with baseline rewards.
Partially-linear bandits. Greenewald et al. (2017) introduced a linear contextual bandit with a time-varying baseline and a TS algorithm with $\tilde{O}(d^{2}\sqrt{T})$ regret, where they used the inverse propensity-weighted observed reward as a pseudo-reward. By explicitly modeling the baseline, we obtain a pseudo-reward with lower variance. Krishnamurthy et al. (2018) improved this to $\tilde{O}(d\sqrt{T})$ regret using a centered RLS estimator, eliminating sub-optimal actions, and choosing a feasible distribution over actions. Kim & Paik (2019) proposed a less restrictive, easier to implement, and faster algorithm with a tight regret upper bound. Our regret bound (see Section 4) involves similar rates but is based on a different asymptotic regime that is not directly comparable due to the presence of an increasing pool of individuals.
Nonlinear bandits. (Li et al., 2017; Wang et al., 2019; Kveton et al., 2020) discussed generalized linear contextual bandit algorithms that accommodate nonlinear relationships via parametric link functions in a similar fashion to generalized linear models Nelder & Wedderburn (1972); McCullagh (2019). Other work (e.g., Snoek et al. (2015); Riquelme et al. (2018); Zhang et al. (2019); Wang & Zhou (2020)) allowed non-parametric relationships in both the baseline reward model and advantage function via deep neural networks; however, these approaches typically lack strong theoretical guarantees and are not designed for longitudinal settings in which pooling offers substantial benefit.
Graph bandits. In the study conducted by Cesa-Bianchi et al. (2013), individual-specific linear models were employed, accompanied by a combinatorial Laplacian penalty to encourage similarity among users’ learned models. This approach yielded a regret bound of $\tilde{O}(nd\sqrt{T})$. Building upon this work, Yang et al. (2020) made further improvements by utilizing a penalty involving the random walk graph Laplacian. Their approach offers the following benefits: (1) it achieves a regret bound of $\tilde{O}(\Psi d\sqrt{nT})$ for some $\Psi\in(0,1)$ and (2) it reduces computational complexity from quadratic to linear by utilizing a first-order approximation to matrix inversion.
Double Machine Learning. Chernozhukov et al. (2018) introduced the DML framework, which provides a general approach to obtain $\sqrt{n}$-consistency for a low-dimensional parameter of interest in the presence of a high-dimensional or “highly complex” nuisance parameter. This framework combines Neyman orthogonality and cross-fitting techniques, ensuring that the estimator is insensitive to the regularization bias produced by the machine learning model. Moreover, it allows us to stay agnostic towards the specific machine learning algorithm while considering the asymptotic properties of the estimator. Later, a number of meta-learner algorithms were developed to leverage the DML framework and provide more precise and robust estimators (Hill, 2011; Semenova & Chernozhukov, 2021; Künzel et al., 2019; Nie & Wager, 2021; Kennedy, 2020).
Doubly Robust Bandits. Kim et al. (2021; 2023) use doubly robust estimators for contextual bandits in both the linear and generalized linear settings, respectively. They use them to obtain a novel regret bound with improved dependence on the dimensionality. In our setting, we use a doubly robust pseudo-reward (robust to either the propensity weights or the mean reward estimate being incorrect) in order to debias explicitly modeling the mean reward. We leave combining our approach with theirs for improved dependence on the dimensionality as future work.
3 Model and Problem Statement
@@
We consider a doubly-indexed contextual bandit with a control action ($a=0$) and $K$ non-baseline arms corresponding to different actions or treatments. Individuals $i=1,2,\ldots$ enter sequentially with each individual observed at a sequence of decision points $t=1,2,\ldots$. For each individual $i$ at time $t$, a context vector $S_{i,t}\in\mathcal{S}$ is observed, an action $A_{i,t}\in\{0,\ldots,K\}:=[K]$ is chosen, and a reward $R_{i,t}\in\mathbb{R}$ is observed. In this paper, we assume the conditional model for the observed reward given state and action, i.e., $\mathbb{E}\left[R_{i,t}|S_{i,t}=s,A_{i,t}=a\right]:=r_{i,t}(s,a)$, is given by
$\begin{split}r_{i,t}(s,a)&=\boldsymbol{x}(s,a)^{\top}\theta_{i,t}\delta_{a>0}+g_{t}(s),\end{split}$ (1)
where $\boldsymbol{x}(s,a)\in\mathbb{R}^{p\times 1}$ is a vector of features of the state and action, $\delta_{a>0}$ is an indicator that takes value $1$ if $a>0$ and 00 otherwise, and $g_{t}(s)$ is a baseline reward that is observed when individuals are randomized to not receive any treatment. This can be an arbitrary, potentially nonlinear function of state $s$ and time $t$. Equation (1) is equivalent to assuming a linear differential reward for any $a>0$; i.e., $\Delta_{i,t}(s,a):=r_{i,t}(s,a)-r_{i,t}(s,0)$ is linear in $\boldsymbol{x}(s,a)$, whose parameter $\theta_{i,t}\in\mathbb{R}^{p}$ is allowed to depend both on the individual $i$ and time $t$.
To mimic real-world recruitment where individuals may not enter a study all at once, we consider a study that proceeds in stages. Figure 3 in Appendix A visualizes this sequential recruitment. At stage $1$, the first individual is recruited and observed at time $t=1$. At stage $k$, individuals $j\leq k$ have been observed for $k-j+1$ decision times respectively. Then each individual $j\in[k+1]$ is observed in a random order at their next time step. Let $\mathcal{H}_{i,t}$ denote the observation history up to time $t$ for individual $i$.
We make the following two standard assumptions as in Abeille & Lazaric (2017).
@@
The reward is observed with additive error $\epsilon_{i,t}$, conditionally mean 0 (i.e., $\mathbb{E}[\epsilon_{i,t}|\mathcal{H}_{i,t}]=0$) sub-Gaussian with variance $\sigma^{2}$: $\mathbb{E}[\exp(\eta\epsilon_{i,t})|\mathcal{H}_{i,t}]\leq\exp(\eta^{2}\sigma^{2}/2)$ for $\eta>0$.
@@
We assume $\|\boldsymbol{x}(s,a)\|\leq 1$ for all contexts and actions and that there exists $B\in\mathbb{R}^{+}$ such that $\|\theta_{i,t}\|\leq B$, $\forall i,t$ and $|g_{t}(s)|\leq B\ \forall s,t$ and $B$ is known.
Here we consider stochastic policies $\pi_{i,t}:\mathcal{H}_{i,t}\times\mathcal{S}\to\mathcal{P}([K])$, which map the observed history $\mathcal{H}_{i,t}$ and current context to a distribution over actions $[K]$. Let $\pi_{i,t}(a|s)$ denote the probability of action $a\in[K]$ given current context $s\in\mathcal{S}$ induced by the map $\pi_{i,t}$ for a fixed (implicit) history.
3.1 DML and Doubly Robust Differential Reward
We first consider a single individual $i$ under a time-invariant linear differential reward, so that $\theta_{i,t}=\theta\in\mathbb{R}^{p}$. If the differential reward $\Delta(s_{i,t},a_{i,t})$ was observed, we could apply ridge regression with a linear model of the form $\boldsymbol{x}(s_{i,t},a_{i,t})^{\top}\theta$ and a ridge penalty of $\lambda\|\theta\|_{2}^{2}$. However, the differential reward is unobserved: we instead consider an inverse-probability weighted (IPW) estimator of the differential reward based on the available data:
$\begin{split}&\mathbb{E}\left[\left(\frac{\delta_{A_{i,t}=\bar{a}}}{1-\pi_{i,t}(0|s)}-\frac{\delta_{A_{i,t}=0}}{\pi_{i,t}(0|s)}\right)R_{i,t}|s_{i,t},\bar{a}_{i,t}\right]=\Delta_{i,t}(s_{i,t},\bar{a}_{i,t})\end{split}$ (2)
where $\bar{a}_{i,t}\in[K]$ denotes the potential non-baseline arm that may be chosen if the baseline arm is not chosen; i.e., randomization is restricted to be between $A_{i,t}=\bar{a}_{i,t}$ and 00. Given the probabilities in the denominators are known, the estimator is unbiased and therefore can replace the observed reward in the Thompson sampling framework.
We refer to this $\Delta_{i,t}(s_{i,t},\bar{a}_{i,t})$ as the differential reward. Below, we define a pseudo-reward with the same expectation in reference to pseudo-outcomes from the causal inference literature (Bang & Robins, 2005; Kennedy, 2020). Let $f_{i,t}(s,a)$ be a working model for the true conditional mean $r_{i,t}(s,a)$. Then, following connections to pseudo-outcomes and doubly-robust (DR) estimators (Kennedy, 2020; Shi & Dempsey, 2023), we define the pseudo-reward $\tilde{R}_{i,t}^{f}(s,\bar{a})$ given state $S_{i,t}=s$ and potential arm $\bar{a}$
$\tilde{R}_{i,t}^{f}(s,\bar{a})\equiv\frac{\left(R_{i,t}-f_{i,t}(s,A_{i,t})\right)}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}+\Delta^{f}_{i,t}(s,\bar{a})$ (3)
@@ 3.1 DML and Doubly Robust Differential Reward
After exploring the properties of the pseudo-reward, an important question arises regarding how we can learn the function $f(s,a)$ using observed data. We hereby provide two options, each based on different assumptions. Option 1 utilizes supervised learning methods and cross-fitting to accurately learn the function while avoiding overfitting as demonstrated in Chernozhukov et al. (2018) and Kennedy (2020). Our model (2) admits sample splitting across time under the assumption of additive i.i.d. errors and no delayed or spill-over effects. Such an assumption is plausible in the mHealth setting where we do not expect an adversarial environment.
In the following, we explain sample splitting as a function of time $t$ as we currently consider a single individual $i$. Step 1: Randomly assign each time $t$ to one of $M$-folds. Let $I_{m}(t)\subseteq\{1,\ldots,t\}$ denote the $m$-th fold as assigned up to time $t$ and $I_{m}^{\complement}(t)$ denote its complement. Step 2: For each fold at each time $t$, use any supervised learning algorithm to estimate the working model for $r_{i,t}(s,a)$ denoted $\hat{f}_{i,t}^{(m)}(s,a)$ using $I_{m}^{\complement}$. Step 3: Construct the pseudo-outcomes using (3) and perform weighted, penalized regression estimation by minimizing the loss function:
$\begin{split}\sum_{m=1}^{M}\sum_{t\in I_{m}(T)}\tilde{\sigma}_{i,t}^{2}&\bigg{(}\tilde{R}_{i,t}^{\hat{f}^{(m)}}-\boldsymbol{x}(s_{i,t},a_{i,t})^{\top}\theta\bigg{)}^{2}\end{split}$ (4)
with ridge penalty $\lambda\|\theta\|_{2}^{2}$, where $\tilde{\sigma}_{i,t}^{2}=\pi_{i,t}(0|s_{i,t})\cdot(1-\pi_{i,t}(0|s_{i,t}))$. The weights are a consequence of unequal variances due to the use of DR estimators; i.e., $\text{var}(\tilde{R}_{i,t}^{f})$ is inversely proportional to $(\tilde{\sigma}_{i,t}^{2})^{2}$.
We explore an alternative, Option 2, based on recent work that avoids sample splitting via the use of stable estimators Chen et al. (2022). To relax the i.i.d. error assumption to Assumption 1, we only update $f_{i,t}(s,a)$ using observed history data in an online fashion, fixing pseudo-outcomes at each stage based on the current estimate of the nonlinear baseline. See Appendix C for further discussion.
Finally, in order to obtain guarantees for this DML approach, we make the following two assumptions: the first is on the convergence in $L^{2}$ (with expectation over states and actions) of our estimate $f$ to the true mean reward and boundedness of the estimator. Similar assumptions were made in Chen et al. (2022). The second is that the weights are bounded below, which would be a consequence if the probability of taking no action is bounded above and below, an assumption made in Greenewald et al. (2017) .

### v3

# A Robust Mixed-Effects Bandit Algorithm for Assessing Mobile Health Interventions

@@ Abstract
Mobile health leverages personalized, contextually-tailored interventions optimized through bandit and reinforcement learning algorithms. Despite its promise, challenges like participant heterogeneity, nonstationarity, and nonlinearity in rewards hinder algorithm performance. We propose a robust contextual bandit algorithm, termed “DML-TS-NNR”, that simultaneously addresses these challenges via (1) modeling the differential reward with user- and time-specific incidental parameters, (2) network cohesion penalties, and (3) debiased machine learning for flexible estimation of baseline rewards. We establish a high-probability regret bound that depends solely on the dimension of the differential reward model. This feature enables us to achieve robust regret bounds even when the baseline reward is highly complex. We demonstrate the superior performance of the DML-TS-NNR algorithm in a simulation and two off-policy evaluation studies.
@@ 1 Introduction
Mobile health (mHealth) uses smart devices to deliver digital notification interventions to users. These notifications nudge users toward healthier attitudes and behaviors. Because mHealth applications can monitor and react to users and their environment, they offer the promise of personalized, contextually-tailored interventions. In practice, achieving this promise requires bandit or reinforcement learning (RL) algorithms that can accurately learn mHealth intervention effects, including how effects vary by user, over time, and based on context. In this regard, contextual bandit algorithms are appealing due to strong empirical performance in other settings, their ability to customize intervention decisions based on changing contexts, and their simplicity and extensibility relative to full RL algorithms.
One unique aspect of mHealth (Greenewald et al., 2017) is the presence of a “do-nothing” action in which the mHealth app does not intervene. As mHealth rewards typically involve human decision making (e.g., whether the user chooses to exercise), the distribution of the the corresponding baseline reward—the reward under the “do-nothing” action—typically evolves in a complex, nonstationary fashion as time progresses and contexts change. In contrast, the treatment effects—the difference in the expected value of the reward relative to the “do-nothing” action—tend to be much more stable and can be adequately modeled using classical stationary models; e.g., linear models (Robinson, 1988).
Greenewald et al. (2017) introduced an action-centered contextual bandit algorithm with sub-linear regret in this setting by replacing the observed reward, $R_{t}$, with a “pseudo-reward,” $\tilde{R}_{t}$, where $t$ indexes the time points. Denoting the binary action as $A_{t}\in\{0,1\}$ with $A_{t}=0$ corresponding to the “do-nothing” action, the pseudo-reward for $A_{t}$ can be written as
$\tilde{R}_{t}:=(A_{t}-\pi_{t})R_{t}=\pi_{t}(1-\pi_{t})R_{t}\left(\frac{A_{t}}{\pi_{t}}-\frac{1-A_{t}}{1-\pi_{t}}\right),$
where $\pi_{t}$ is the (known) probability of treatment assignment from the Thompson-sampling algorithm. We took an additional step to derive an equivalent expression (second equality), which reveals that $\tilde{R}_{t}$ is proportional to an inverse-probability-weighted (IPW) estimator. Crucially, $\tilde{R}_{t}$ provides an unbiased estimate of the treatment effect regardless of the non-stationarity of the baseline outcome.
Although Greenewald et al. (2017)’s analysis shows sub-linear regret in nonstationary settings, our simulations show that it is outperformed by other methods over finite time horizons. The reasons include its failure to address treatment effect heterogeneity, pool information across users, and model the baseline outcome, which leads to highly variable pseudo-rewards and slow learning. We propose to generalize their method to overcome these limitations by
  1. 1.
    imposing hierarchical structure in the treatment effect model with shared structural parameters and “incidental parameters” for users and time,
  2. 2.
    efficiently pooling across users and time via nearest-neighbor regularization (NNR),
  3. 3.
    de-noising rewards with flexible supervised learning models via the de-biased machine learning (DML) framework of Chernozhukov et al. (2018).
We name the resulting method DML-TS-NNR to highlight its usage of DML, NNR, and Thompson Sampling. The primary contributions are (a) the DML-TS-NNR method, (b) a high-probability regret bound that relies solely on the dimension of the differential reward model, which is typically much smaller than that of the complex baseline reward model, and (c) empirical comparisons demonstrating the superior performance of DML-TS-NNR in simulation and two real-world mHealth studies.
@@ 2 Related Work
Related work in statistics considers treatment effect estimation with longitudinal data. Cho et al. (2017) present a semiparametric random effects method for estimating subject-specific static treatment effects. Qian et al. (2020) discusses the statistical challenges of applying linear mixed-effects models to mHealth data and shows that the resulting estimates may lack a causal interpretation.
In the bandit literature, several works address a growing pool of users. Ma et al. (2011) introduced NNR in recommender systems using homophily principles—similar nodes are more likely to be connected than dissimilar ones (McPherson et al., 2001). Cesa-Bianchi et al. (2013) adapted NNR for bandit settings, improving regret. Subsequent improvements focused on scalability and algorithm modifications for stronger regret bounds (Vaswani et al., 2017; Yang et al., 2020).
Other bandit approaches explicitly address the longitudinal setting in which treatment effects may evolve over time. The intelligentpooling method of Tomkins et al. (2021) employs a Gaussian linear mixed-effects model with both user- and time-specific random effects. Hu et al. (2021) provides a related approach based on generalized linear mixed effects models. Unlike our approach, both methods require the (generalized) linear conditional mean model to be correctly specified.
Aouali et al. (2023) and Lee et al. (2024) also propose bandit algorithms with mixed effects; methodologically, however, their use of mixed effects is dissimilar from that of the works above and our own. Aouali et al. (2023) employs random effects to relate treatments with categories (e.g., movies within a genre), and Lee et al. (2024) considers the case in which a single, discrete action leads to a multivariate reward that is affected by a user-specific random effect. In contrast, the other approaches listed above (and our own) do not assume any relationships among the actions, and they associate each discrete action with a single scalar-valued reward in an iterative fashion.
Other work develops semiparametric methods that lead to sublinear regret without requiring correct specification of the conditional mean reward. Krishnamurthy et al. (2018) and Kim & Paik (2019) propose improvements to the approach of Greenewald et al. (2017) that lead to stronger regret bounds. Kim et al. (2021, 2023) develop a doubly-robust approach for both linear and generalized linear contextual bandits. However, in contrast to Tomkins et al. (2021), Hu et al. (2021), and our approach, these semiparametric methods are not immediately applicable to longitudinal settings.
3 Setting
3.1 Problem Statement
@@
Recognizing the connection between mHealth policy learning and contextual bandit methods, we now consider a contextual multi-armed bandit environment with one control arm ($a=0$) and $q$ non-baseline arms corresponding to different actions or treatments. Individuals are indexed by $i=1,2,\ldots$, and decision points are indexed by $t=1,2,\ldots$. For each individual $i$ at time $t$, the algorithm observes a context vector $S_{i,t}\in\mathcal{S}$, chooses an action $A_{i,t}\in[q]:=\{0,\ldots,q\}$, and receives a reward $R_{i,t}\in\mathbb{R}$. The differential reward describes the difference in expected rewards between choosing a non-baseline arm and the control arm, defined as $\Delta_{i,t}(s,a):=\mathbb{E}\left[R_{i,t}|S_{i,t}=s,A_{i,t}=a\right]-\mathbb{E}\left[R_{i,t}|S_{i,t}=s,A_{i,t}=0\right]$.
We denote the observation history up to time $t$ for individual $i$ by $\mathcal{H}_{i,t}:=(S_{i1},A_{i1},R_{i1},\dots,R_{it})$. Actions are selected according to stochastic policies, $\pi_{i,t}:\mathcal{H}_{i,t}\times\mathcal{S}\to\mathcal{P}([q])$. We use $\pi_{i,t}(a|s)$ to denote the probability of action $a\in[q]$ given current context $s\in\mathcal{S}$ for a fixed (implicit) history. As in Greenewald et al. (2017), we assume that the probability of control action is bounded:
@@
There exists $10<\pi_{\min},\pi_{\max}<$ 1 s.t $\pi_{\min}<\pi_{i,t}(0|a),1-\pi_{i,t}(0|a)<\pi_{\max}\forall i,t,a$.
As Greenewald et al. (2017) noted, this assignment probability controls the number of messages that participants receive. It ensures that participants receive sufficient messages that they do not disengage, but not so many that they are overwhelmed or fatigued. We now define
$A^{*}=\operatorname{argmax}_{a\in[q]}\mathbb{E}\left[R_{i,t}|S_{i,t},A_{i,t}=a\right],\quad\bar{A}^{*}=\operatorname{argmax}_{a\in[q]\backslash\{0\}}\Delta_{i,t}(S_{i,t},a),$
the optimal treatment arm (including control) and the alternative arm, respectively. Respecting the restrictions in Assumption 1, we can now express the optimal policy, $\pi_{i,t}^{*}$, in two cases:
  1. 1.
    $A^{*}=0$: $\pi_{i,t}^{*}(0|S_{i,t})=\pi_{\max}$ and $\pi_{i,t}^{*}(\bar{A}^{*}|S_{i,t})=1-\pi_{\max}$.
  2. 2.
    $A^{*}>0$: $\pi_{i,t}^{*}(A^{*}|S_{i,t})=1-\pi_{\min}$ and $\pi_{i,t}^{*}(0|S_{i,t})=\pi_{\min}$.
These cases represent achieving the appropriate boundary condition given the optimal action, i.e., maximum chance of “do-nothing” action if $A^{\star}=0$ and minimum chance if $A^{\star}\not=0$. Otherwise, $\pi_{i,t}^{*}(a|S_{i,t})=0$. In Section 5.2, we define regret relative to this optimal policy, $\pi_{i,t}^{*}$.
Lastly, we consider a study design that progresses in stages, where individuals enter the study sequentially (Friedman et al., 2010), as illustrated in Figure 3 in Appendix A. Staged recruitment serves two purposes in our analysis: it increases the fidelity of the theoretical setting by mimicking real-world recruitment strategies, and it allows us to estimate changes in the differential reward over time. Concretely, we first observe individual $i=1$ at time $t=1$. Then we observe individuals $i=(1,2)$ at times $t=(2,1)$. Subsequent stages $k$ involve individuals $i\leq k$, each having had $k-i+1$ observations, respectively. Define $\mathcal{O}_{k}=\{(i,t):i\leq k\>\&\>t\leq k+1-i\}$ as the double index set of all observations in stage $k$. A notation summary is presented in Appendix G.
3.2 Doubly Robust Differential Reward
Following Greenewald et al. (2017), we assume that the differential reward is linear:
$\Delta_{i,t}(s,a)=x(s,a)^{\top}\theta_{i,t},$ (1)
where $x(s,a)\in\mathbb{R}^{p\times 1}$ is a feature vector depending on the context and action. We allow the baseline reward, $g_{t}(s)$, to be a complex, nonlinear, and nonstationary function of context and time. Combined, these two assumptions imply a partially linear model for the conditional mean reward:
$\begin{split}\mathbb{E}\left[R_{i,t}|S_{i,t}=s,A_{i,t}=a\right]&=x(s,a)^{\top}\theta_{i,t}\delta_{a>0}+g_{t}(s),\end{split}$ (2)
where $\delta_{a>0}$ is an indicator of a non-baseline action. By incorporating both the linear predictor $x(s,a)^{\top}\theta_{i,t}\delta_{a>0}$ and the nonlinear baseline $g_{t}(s)$, we can maintain interpretability of the action effects while also allowing flexibility in specifying the observed rewards. The individual rewards are then realized according to the model
$R_{i,t}=x(s,a)^{\top}\theta_{i,t}\delta_{a>0}+g_{t}(s)+\epsilon_{i,t},$
where we make the following standard (Abeille & Lazaric, 2017) assumption regarding $\epsilon_{i,t}$:
@@ 3.2 Doubly Robust Differential Reward
$\epsilon_{i,t}$ is conditionally mean zero (i.e., $\mathbb{E}[\epsilon_{i,t}|\mathcal{H}_{i,t}]=0$) and sub-Gaussian with variance $\sigma^{2}$: $\mathbb{E}[\exp(\eta\epsilon_{i,t})|\mathcal{H}_{i,t}]\leq\exp(\eta^{2}\sigma^{2}/2)$ for $\eta>0$.
Under this model assumption, the IPW estimator of Greenewald et al. (2017) produces an unbiased estimate of $\Delta_{i,t}(s,a)$; however, it suffers from high variance, which leads to suboptimal performance in practice. Our proposed algorithm DML-TS-NNR, however, relies on an improved estimator, $\tilde{R}_{i,t}^{f}(s,\bar{a})$, that is also unbiased but achieves lower variance by denoising the rewards according to a working model, $f_{i,t}(s,a)$, for the conditional mean $r_{i,t}(s,a):=\mathbb{E}\left[R_{i,t}|S_{i,t}=s,A_{i,t}=a\right]$:
@@
where $\bar{a}$ is a pre-selected non-baseline arm; i.e., we condition on $A_{i,t}\in\{0,\bar{a}\}$. The first component is a model-based prediction of the differential reward, and the second component is a debiasing term that uses IPW to correct for potential errors in specifying $f$. This improved estimator is a form of pseudo-outcome, a common technique in the causal inference literature (Bang & Robins, 2005; Kennedy, 2020; Nie & Wager, 2021). We refer to it as a Doubly Robust Differential Reward because it produces an unbiased estimate of $\Delta_{i,t}(s,a)$ as long as either $\pi_{i,t}$ or $f_{i,t}$ is correctly specified (see proof in Appendix F.1); though, in practice, the benefit stems from (a) the robustness to misspecification of $f$ (because $\pi_{i,t}$ is known) and (b) the variance reduction introduced by de-noising the outcomes. Lemma 5 and Remark 2 in Appendix F.2 provide a proof and discussion of the variance reduction. In the following, we abbreviate $\tilde{R}_{i,t}^{f}(s,a)$ as $\tilde{R}_{i,t}^{f}$, with the state and action implied.
4 Algorithm Components
We now construct $\tilde{R}_{i,t}^{f}$, introduce the weighted least-squares loss function used by DML-TS-NNR to estimate parameters, and use nearest neighbor regularization to pool information across users.
4.1 Debiased Machine Learning
To construct $\tilde{R}_{i,t}^{f}$, we fit a working model $f_{i,t}$ for the conditional mean reward via the DML framework, which requires three main conditions: (a) Neyman Orthogonality - making the pseudo-outcome $\tilde{R}_{i,t}^{f}$ robust to misspecification of $f_{i,t}$, (b) sufficiently fast convergence of the working conditional mean model $f_{i,t}$, and (c) independence of $f_{i,t}$ and the observed outcomes. (a) is ensured by the structure of the doubly-robust differential reward. (b) requires us to introduce the following assumption: